# A Systematic Literature Review on Large Language Models for Automated Program Repair

### v3

@@ 8.2. What input forms are software bugs transformed into when utilizing LLMs?
2025 22 (Dai et al., 2025) (Garg et al., 2023) (Ehsani et al., 2025b) (Ma et al., 2025a) (Wei et al., 2025) (Xu et al., 2023) (Xia et al., 2025) (Yuan et al., 2024) (Zhang et al., 2024c) (Nong et al., 2025) (Cao et al., 2023) (Feng et al., 2025) (Hu et al., 2025) (Kong et al., 2025) (Lin et al., 2025) (Ma et al., 2025b) (Parasaram et al., 2025) (Xu et al., 2025) (Yang et al., 2024e) (Huang et al., 2025b) (Zhang et al., 2025) (Silva et al., 2023)
@@ 8.5. How often do the collected LLM-based APR papers provide publicly available artifacts?
9. Challenges & Opportunities of LLM-based APR
Our study reveals the following crucial challenges and practical guidelines for future LLM-based APR work.
APR as a Part of Fully Autonomous Programming. As mentioned in Section 2, existing APR approaches operate under a “near-correct assumption” (Offutt and Lee, 1994; DeMillo et al., 2006), which posits that experienced programmers are capable of writing programs that are almost correct, requiring only minimal modifications to rectify bugs and ensure compliance with all test cases. This assumption has long served as the foundation for APR research development. However, the evolution of LLMs and their application in programming suggests a future where APR can transcend its traditional boundaries, moving towards a more holistic approach in conjunction with fully autonomous programming. In this context, APR can be reimagined not only just as a tool for correcting minor coding errors, but as an integral component of a self-correcting, self-improving system that iteratively enhances the quality of automatically generated code. Fortunately, we have observed initial explorations into combining repair with programming, yet these efforts are far from sufficient. For example, Fan et al. (Fan et al., 2023a) employ LLMs to fix buggy solutions generated by themselves, and recent studies (Zhang et al., 2023k; Xia and Zhang, 2024) iteratively refine auto-generated code with dynamic execution.
In the future, the opportunities presented by integrating APR with fully autonomous programming are vast (Lyu et al., 2024). First, it is flexible to implement collaborative Human-AI Programming tools (Liang et al., 2021), where the initial code is written by developers and continuously optimized and repaired by LLMs. Particularly for complex problem-solving, LLMs suggest innovative solutions that had not been considered by human programmers. This partnership could accelerate development cycles, reduce the burden of debugging on human developers, and lead to more creative and effective solutions. Second, the general knowledge inherent in LLMs endows them with the capability to support multiple downstream tasks, which opens up the possibility of bridging the gap between code generation, testing, debugging, and fixing. For example, fault localization is a precondition of patch generation, while patch validation can reflect the accuracy of fault localization, making these takes interconnected. Thus, it is promising to explore the capabilities of LLMs in these connected tasks using real-time feedback within a unified framework.
More Attention about Repair Costs of LLMs.. As mentioned in Section 6.1, the literature tends to employ the growing size of LLMs to achieve better performance, such as from T5-250M in CIRCLE to Codex-12B in InferFix. This trend is reasonable as Xia et al. (Xia et al., 2023b) demonstrate that larger models typically generate more correct patches for software bugs. However, APR, being a task highly relevant to humans, the millions, billions and even trillions of parameters pose significant challenges in the development workflow. First, training LLMs with billions (or even more) of parameters is a highly time-consuming and resource-intensive process. The GPU resources required for such training are often prohibitively expensive, making them inaccessible for many researchers in both academic and industrial settings. For example, in Section 6.2, we observe that most fine-tuning-LLM-based APR studies utilize CodeT5/T5 or similar-sized models, except for InferFix from Microsoft, which fine-tunes Codex. However, InferFix is proposed by the world-leading technology company Microsoft and trained with industrial-grade hardware, i.e., 64 32-GB-V100-GPUs. Second, despite the increased likelihood of generating correct patches with larger models, the patch generation time cost also increases. For example, Jiang et al. (Jiang et al., 2023a) demonstrate PLBART takes 0.70–0.89 seconds on average to generate a correct patch, CodeGen, although fixing more bugs than PLBART with more parameters, requires 3.64–13.88 seconds on average to generate a correct patch. Third, the increase in patch generation time further compresses the time available for patch validation, as developers need to spend more time waiting for the model to complete its inference. For example, Shi et al. (Shi et al., 2022) demonstrate that the vast storage and runtime memory consumption, coupled with high inference latency, make these LLMs prohibitive for integration in modern IDEs, especially on resource-constrained or real-time terminal devices.
In the future, to address the first challenge, it is promising to explore the potential of parameter-efficient fine-tuning approaches on APR, such as prefix-tuning and low-rank adaptation (Ding et al., 2023). In addition, recent sparsity-aware optimization frameworks such as SparseGPT (Frantar and Alistarh, 2023), SparseML (Magic, 2023), and DeepSparse (Neural Magic, 2024) demonstrate promising directions for improving inference efficiency of LLMs, which could also benefit LLM-based APR in future work. To address the second challenge, researchers can optimize the size of LLMs without significantly compromising their performance, such as model pruning, quantization, and knowledge distillation (Shi et al., 2022). To address the third challenge, we recommend boosting patch validation with advanced strategies, such as mutation testing (DeMillo et al., 2006; Xiao et al., 2024), or utilizing LLMs to rank candidate patches before validation.
Human Study with LLMs. As summarized in Section 7, with the introduction of LLMs, APR has achieved groundbreaking progress in terms of the number of correctly fixed bugs in popular benchmarks, exemplified by the advancements on Defects4J-1.2 from 64 bugs by CIRCLE, to 74 bugs by AlphaRepair, and 82 bugs by GAMMA. The progress prompts us to consider whether LLMs have indeed facilitated improvements in real-world debugging and how LLM-based APR has broader implications for developers’ daily activities. Previous research (Zhang et al., 2022c; Winter et al., 2022) highlights the potential pitfalls of developing APR tools without adequate feedback from developers, which could compromise their effectiveness in real-world deployment. However, there remains a significant gap in our understanding of how software engineers tackle software problems in practical settings, including their use of dedicated debugging tools and their expertise in debugging techniques.
Thus, in the future, researchers should conduct human studies to gain deeper insights into the maturity and reliability of LLM-based APR tools in terms of human factors. Possible directions are to investigate whether LLMs can assist developers in reducing the debugging cost, such as fixing more bugs, accelerating the bug-fixing process, and handling more complex bugs. Besides, it would be valuable to investigate developers’ perceptions and interactions with LLMs based on their practical experiences and established debugging practices.
Exploring More and Rare Repair Scenarions. As summarized in Section 7, we observe that most existing LLM-based APR studies are concentrated on a limited number of bug types, particularly semantic bugs. However, there exist some rare repair scenarios that benefit less from LLMs, such as hardware bugs (one paper) and concurrency bugs (zero paper). The key challenge lies in insufficient training data from which LLMs can learn.
We suggest that future work concentrates on three possible directions to broaden the scope of LLM applications for more repair scenarios, such as software requirement (Xie et al., 2023) and fuzzing (Zhang et al., 2023j). First, transfer learning is an effective training approach for rare scenarios. We can first fine-tune LLMs with abundant data in a source scenario and then utilize a small amount of data to transfer the acquired knowledge to a target scenario. The source and target scenarios should have similar data distributions. For example, Zhang et al. (Zhang et al., 2023d) demonstrate that transferring learning from bug-fixing can improve the vulnerability repair performance of five LLMs by 9.40% on average. Second, it is promising to utilize the cloze-style APR with repair patterns to generate patches for rare scenarios. Second, it is promising to utilize the cloze-style APR with repair patterns to generate patches for rare scenarios. Unlike fine-tuning, which requires substantial labeled data, cloze-style APR effectively leverages expert domain knowledge to guide LLMs in performing APR based on their existing pre-trained knowledge without additional retraining. For example, researchers design repair patterns and exploit LLMs’ pre-training objective (i.e., predicting masked tokens) to directly generate correct patches by filling masked tokens under these patterns. This makes the cloze-style APR particularly effective for scenarios with limited training data or domain-specific bug types. Besides, for scenarios not previously encountered by LLMs, researchers can employ unlabeled data in the target scenario to learn the data distribution of the project or language under test in an unsupervised learning setting. Third, in-context learning and prompt engineering are feasible solutions for directly querying billion-parameter LLMs to generate correct code in a target scenario, given that these models are immensely large and encompass virtually all data available on the Internet.
@@ 9. Challenges & Opportunities of LLM-based APR
Integration with Off-the-Shelf APR. As mentioned in Section 7, researchers typically utilize LLMs as core backbones to design novel repair approaches in an end-to-end setting, such as sequence-to-sequence learning (Yuan et al., 2022; Berabi et al., 2021). Parallelly, the community has seen some explorations treating LLMs as components integrated into existing repair workflow. These studies attempt to boost the capabilities of off-the-shelf APR approaches instead of proposing new techniques. For example, CURE (Jiang et al., 2021) combines GPT and CoCoNut (Lutellier et al., 2020) to capture code syntax for the APR task. Built on top of DLFix (Li et al., 2020), DEAR (Li et al., 2022b) attempts to fix multi-hunk, multi-statement bugs by fine-tuning BERT to learn fixing-together relationships among statements, i.e., identifying if two statements are needed to be fixed together or not. Recently, GAMMA (Zhang et al., 2023e) integrates LLMs into the traditional template-based APR TBar by querying LLMs to generate masked code tokens instead of retrieving donor code from local files. These efforts demonstrate the potential of integrating LLMs with off-the-shelf APR techniques, yet there is currently a lack of more in-depth work in this area.
@@
In the future, researchers could attempt to combine LLMs with more traditional APR techniques. For example, it is promising to utilize LLMs to help SMT solvers generate patches for constraint-based APR, or feed search algorithms with LLM-generated potential patches to build search space for heuristic-based APR. Besides, domain-specific repair techniques can benefit from the powerful code-understanding capabilities of LLMs, thus extending to a broader range of repair scenarios. For example, we can design fix templates for specific scenarios, such as static warnings, and then utilize the general knowledge contained in LLMs to generate correct patches.
@@
Data Leakage Issue. As highlighted in Section 8.1, Zhang et al. (Zhang et al., 2023k) identify that existing repair benchmarks have been inadvertently included in the pre-training data of popular LLMs, such as ChatGPT, through web scraping and other methods. For example, ChatGPT is able to enumerate all projects within Defects4J (Just et al., 2014), one of the most popular APR benchmarks. Researchers (Zhang et al., 2023e) can ascertain the exposure for open-source LLMs by inspecting pre-training data against benchmarks. It is significantly more challenging with more powerful black-box LLMs due to a lack of training details. However, as preliminary explorations, they are mainly limited in the quantity of involved bugs and the variety of bug types. For example, all of them are created from programming problems with only small-scale or medium-scale buggy solutions, without delving into large-scale, real-world projects that encompass complex API calls, such as Defecst4J (Just et al., 2014). More importantly, data leakage issues in other repair scenarios, such as security vulnerabilities and API misuse, continue to be overlooked. There may be overlaps among datasets across different repair scenarios, such as Defects4J, which also serves as a source for a subset of the API misuse dataset (Zhang et al., 2023f). Overall, the risk of data leakage introduces bias into the benchmarking of existing work, necessitating urgent efforts from researchers to mitigate it.
@@
We recommend that future work can be conducted in the following two directions. First, it is crucial to construct a large-scale benchmark free from data leakage that contains real-world projects so as to evaluate the actual fix capabilities of LLMs in a more practical debugging scenario. Commercial closed-source software, real-time updated programming websites, or manually written programs may serve as potential data sources. Second, considering the variety of bug types that LLMs have been applied to, researchers need to consider and attempt to address the data leakage risk when conducting related studies.
@@ Acknowledgments
This work is supported partially by the National Natural Science Foundation of China (61932012, 62141215, 62372228).

### v4

@@ 8.2. What input forms are software bugs transformed into when utilizing LLMs?
2025 22 (Dai et al., 2025) (Garg et al., 2023) (Ehsani et al., 2025b) (Ma et al., 2025a) (Wei et al., 2025) (Xu et al., 2023) (Xia et al., 2025a) (Yuan et al., 2024) (Zhang et al., 2024c) (Nong et al., 2025) (Cao et al., 2023) (Feng et al., 2025) (Hu et al., 2025) (Kong et al., 2025) (Lin et al., 2025) (Ma et al., 2025b) (Parasaram et al., 2025) (Xu et al., 2025) (Yang et al., 2024e) (Huang et al., 2025b) (Zhang et al., 2025d) (Silva et al., 2023)
@@ 8.5. How often do the collected LLM-based APR papers provide publicly available artifacts?
9. Research Agenda for LLM-based APR
Figure 5. A three-stage future research roadmap for LLM-based APR.
Based on our systematic analysis, this section presents a structured research agenda that outlines concrete and actionable directions for future studies on LLM-based APR. As illustrated in Figure 5, we organize future research directions into three progressive stages: (1) advancing LLM-based APR techniques, (2) rigorous evaluation and benchmarking, and (3) real-world deployment and adoption.
9.1. Stage 1: Advancing LLM-based APR Techniques
9.1.1. Off-the-Shelf APR Integration
@@
Current state. As mentioned in Section 7, researchers typically utilize LLMs as core backbones to design novel repair approaches in an end-to-end setting, such as sequence-to-sequence learning (Yuan et al., 2022; Berabi et al., 2021). Parallelly, the community has seen some explorations treating LLMs as components integrated into existing repair workflow. These studies attempt to boost the capabilities of off-the-shelf APR approaches instead of proposing new techniques. For example, CURE (Jiang et al., 2021) combines GPT and CoCoNut (Lutellier et al., 2020) to capture code syntax for the APR task. Built on top of DLFix (Li et al., 2020), DEAR (Li et al., 2022b) attempts to fix multi-hunk, multi-statement bugs by fine-tuning BERT to learn fixing-together relationships among statements, i.e., identifying if two statements are needed to be fixed together or not. Recently, GAMMA (Zhang et al., 2023e) integrates LLMs into the traditional template-based APR TBar by querying LLMs to generate masked code tokens instead of retrieving donor code from local files. These efforts demonstrate the potential of integrating LLMs with off-the-shelf APR techniques, yet there is currently a lack of more in-depth work in this area.
@@
Opportunities. In the future, researchers could attempt to combine LLMs with more traditional APR techniques. For example, it is promising to utilize LLMs to help SMT solvers generate patches for constraint-based APR, or feed search algorithms with LLM-generated potential patches to build search space for heuristic-based APR. Besides, domain-specific repair techniques can benefit from the powerful code-understanding capabilities of LLMs, thus extending to a broader range of repair scenarios. For example, we can design fix templates for specific scenarios, such as static warnings, and then utilize the general knowledge contained in LLMs to generate correct patches.
9.1.2. Under-explored Repair Scenarios
Current state. As summarized in Section 7, we observe that most existing LLM-based APR studies are concentrated on a limited number of bug types, particularly semantic bugs. However, there exist some rare repair scenarios that benefit less from LLMs, such as hardware bugs (one paper) and concurrency bugs (zero paper). A key challenge lies in the scarcity of labeled training data and the increased complexity associated with these scenarios.
Opportunities. We suggest that future work concentrate on three possible directions to broaden the scope of LLM applications for more repair scenarios, such as software requirement (Xie et al., 2023) and fuzzing (Zhang et al., 2023j).
  • Transfer learning is an effective training approach for rare scenarios. We can first fine-tune LLMs with abundant data in a source scenario and then utilize a small amount of data to transfer the acquired knowledge to a target scenario. The source and target scenarios should have similar data distributions. For example, Zhang et al. (Zhang et al., 2023d) demonstrate that transferring learning from bug-fixing can improve the vulnerability repair performance of five LLMs by 9.40% on average.
  • It is promising to utilize the cloze-style APR with repair patterns to generate patches for rare scenarios. Unlike fine-tuning, which requires substantial labeled data, cloze-style APR effectively leverages expert domain knowledge to guide LLMs in performing APR based on their existing pre-trained knowledge without additional retraining. For example, researchers design repair patterns and exploit LLMs’ pre-training objective (i.e., predicting masked tokens) to directly generate correct patches by filling masked tokens under these patterns. This makes the cloze-style APR particularly effective for scenarios with limited training data or domain-specific bug types. Besides, for scenarios not previously encountered by LLMs, researchers can employ unlabeled data in the target scenario to learn the data distribution of the project or language under test in an unsupervised learning setting.
  • In-context learning and prompt engineering are feasible solutions for directly querying billion-parameter LLMs to generate correct code in a target scenario, given that these models are immensely large and encompass virtually all data available on the Internet.
9.2. Stage 2: Rigorous Evaluation and Benchmarking
9.2.1. Data Leakage Mitigation
@@
Current state. As highlighted in Section 8.1, Zhang et al. (Zhang et al., 2023k) identify that existing repair benchmarks have been inadvertently included in the pre-training data of popular LLMs, such as ChatGPT, through web scraping and other methods. For example, ChatGPT is able to enumerate all projects within Defects4J (Just et al., 2014), one of the most popular APR benchmarks. Researchers (Zhang et al., 2023e) can ascertain the exposure for open-source LLMs by inspecting pre-training data against benchmarks. It is significantly more challenging with more powerful black-box LLMs due to a lack of training details. However, as preliminary explorations, they are mainly limited in the quantity of involved bugs and the variety of bug types. For example, all of them are created from programming problems with only small-scale or medium-scale buggy solutions, without delving into large-scale, real-world projects that encompass complex API calls, such as Defects4J (Just et al., 2014). More importantly, data leakage issues in other repair scenarios, such as security vulnerabilities and API misuse, continue to be overlooked. There may be overlaps among datasets across different repair scenarios, such as Defects4J, which also serves as a source for a subset of the API misuse dataset (Zhang et al., 2023f). Overall, the risk of data leakage introduces bias into the benchmarking of existing work, necessitating urgent efforts from researchers to mitigate it.
@@
Opportunities. We recommend that future work be conducted in the following two directions. First, it is crucial to construct a large-scale benchmark free from data leakage that contains real-world projects so as to evaluate the actual fix capabilities of LLMs in a more practical debugging scenario. Commercial closed-source software, real-time updated programming websites, or manually written programs may serve as potential data sources. Second, considering the variety of bug types that LLMs have been applied to, researchers need to consider and attempt to address the data leakage risk when conducting related studies.
9.2.2. Cost-aware Evaluation
Current state. As mentioned in Section 6.1, the literature tends to employ the growing size of LLMs to achieve better performance, such as from T5-250M in CIRCLE to Codex-12B in InferFix. This trend is reasonable as Xia et al. (Xia et al., 2023b) demonstrate that larger models typically generate more correct patches for software bugs. However, APR, being a task highly relevant to humans, the millions, billions and even trillions of parameters pose significant challenges in the development workflow. First, training LLMs with billions (or even more) of parameters is a highly time-consuming and resource-intensive process. The GPU resources required for such training are often prohibitively expensive, making them inaccessible for many researchers in both academic and industrial settings. For example, in Section 6.2, we observe that most fine-tuning-LLM-based APR studies utilize CodeT5/T5 or similar-sized models, except for InferFix from Microsoft, which fine-tunes Codex. However, InferFix is proposed by the world-leading technology company Microsoft and trained with industrial-grade hardware, i.e., 64 32-GB-V100-GPUs. Second, despite the increased likelihood of generating correct patches with larger models, the patch generation time cost also increases. For example, Jiang et al. (Jiang et al., 2023a) demonstrate PLBART takes 0.70–0.89 seconds on average to generate a correct patch, CodeGen, although fixing more bugs than PLBART with more parameters, requires 3.64–13.88 seconds on average to generate a correct patch. Third, the increase in patch generation time further compresses the time available for patch validation, as developers need to spend more time waiting for the model to complete its inference. For example, Shi et al. (Shi et al., 2022) demonstrate that the vast storage and runtime memory consumption, coupled with high inference latency, make these LLMs prohibitive for integration in modern IDEs, especially on resource-constrained or real-time terminal devices.
Opportunities. In the future, to address the first challenge, it is promising to explore the potential of parameter-efficient fine-tuning approaches on APR, such as prefix-tuning and low-rank adaptation (Ding et al., 2023). In addition, recent sparsity-aware optimization frameworks such as SparseGPT (Frantar and Alistarh, 2023), SparseML (Magic, 2023), and DeepSparse (Neural Magic, 2024) demonstrate promising directions for improving inference efficiency of LLMs, which could also benefit LLM-based APR in future work. To address the second challenge, researchers can optimize the size of LLMs without significantly compromising their performance, such as model pruning, quantization, and knowledge distillation (Shi et al., 2022). To address the third challenge, we recommend boosting patch validation with advanced strategies, such as mutation testing (DeMillo et al., 2006; Xiao et al., 2024), or utilizing LLMs to rank candidate patches before validation.
9.3. Stage 3: Real-world Deployment
9.3.1. Human-centered Adoption
Current state. As summarized in Section 7, with the introduction of LLMs, APR has achieved groundbreaking progress in terms of the number of correctly fixed bugs in popular benchmarks, exemplified by the advancements on Defects4J-1.2 from 64 bugs by CIRCLE, to 74 bugs by AlphaRepair, and 82 bugs by GAMMA. The progress prompts us to consider whether LLMs have indeed facilitated improvements in real-world debugging and how LLM-based APR has broader implications for developers’ daily activities. Previous research (Zhang et al., 2022c; Winter et al., 2022) highlights the potential pitfalls of developing APR tools without adequate feedback from developers, which could compromise their effectiveness in real-world deployment. However, there remains a significant gap in our understanding of how software engineers tackle software problems in practical settings, including their use of dedicated debugging tools and their expertise in debugging techniques.
Opportunities. Researchers should conduct human studies to gain deeper insights into the maturity and reliability of LLM-based APR tools in terms of human factors. Possible directions are to investigate whether LLMs can assist developers in reducing the debugging cost, such as fixing more bugs, accelerating the bug-fixing process, and handling more complex bugs. Besides, it would be valuable to investigate developers’ perceptions and interactions with LLMs based on their practical experiences and established debugging practices.
9.3.2. Programming Workflows
Current state. As mentioned in Section 2, existing APR approaches operate under a “near-correct assumption” (Offutt and Lee, 1994; DeMillo et al., 2006), which posits that experienced programmers are capable of writing programs that are almost correct, requiring only minimal modifications to rectify bugs and ensure compliance with all test cases. This assumption has long served as the foundation for APR research development. However, the evolution of LLMs and their application in programming suggests a future where APR can transcend its traditional boundaries, moving towards a more holistic approach in conjunction with fully autonomous programming. In this context, APR can be reimagined not only just as a tool for correcting minor coding errors, but as an integral component of a self-correcting, self-improving system that iteratively enhances the quality of automatically generated code. Fortunately, we have observed initial explorations into combining repair with programming, yet these efforts are far from sufficient. For example, Fan et al. (Fan et al., 2023a) employ LLMs to fix buggy solutions generated by themselves, and recent studies (Zhang et al., 2023k; Xia and Zhang, 2024) iteratively refine auto-generated code with dynamic execution.
Opportunities. In the future, the opportunities presented by integrating APR with fully autonomous programming are vast (Lyu et al., 2024). First, it is flexible to implement collaborative Human-AI Programming tools (Liang et al., 2021), where the initial code is written by developers and continuously optimized and repaired by LLMs. Particularly for complex problem-solving, LLMs suggest innovative solutions that had not been considered by human programmers. This partnership could accelerate development cycles, reduce the burden of debugging on human developers, and lead to more creative and effective solutions. Second, the general knowledge inherent in LLMs endows them with the capability to support multiple downstream tasks, which opens up the possibility of bridging the gap between code generation, testing, debugging, and fixing. For example, fault localization is a precondition of patch generation, while patch validation can reflect the accuracy of fault localization, making these takes interconnected. Thus, it is promising to explore the capabilities of LLMs in these connected tasks using real-time feedback within a unified framework.
@@ Acknowledgments
This research was supported in part by the Australian Research Council Discovery Projects under Grant No. DP200102491 and DP230101790, the National Natural Science Foundation of China (61932012, 62141215, 62372228), Natural Science Foundation of Jiangsu Province (BK20251458), and Fundamental Research Funds for the Central Universities (AE89991/463).